# Sample-Based Conservative Bias Linear Power Flow Approximations

### v1

@@ Abstract
The power flow equations are central to many problems in power system planning, analysis, and control. However, their inherent non-linearity and non-convexity present substantial challenges during problem-solving processes, especially for optimization problems. Accordingly, linear approximations are commonly employed to streamline computations, although this can often entail compromises in accuracy and feasibility. This paper proposes an approach termed Conservative Bias Linear Approximations (CBLA) for addressing these limitations. By minimizing approximation errors across a specified operating range while incorporating conservativeness (over- or under-estimating quantities of interest), CBLA strikes a balance between accuracy and tractability by maintaining linear constraints. By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy. We illustrate the effectiveness of our proposed approach through several test cases.
@@
Conservative bias linear approximation; power flow approximation
@@ I Introduction
The power flow equations play a central role in the operation and analysis of electrical power systems. These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7]. However, the nonlinearity of the power flow equations induces non-convexities in these problems that pose significant computational challenges.
To address these challenges, researchers have developed various linear approximations such as DC power flow [8], LinDistFlow [9], first-order Taylor expansions of the power flow equations, and other approximations [10]. These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems. However, these linearizations often depend on broad assumptions such as maintaining voltages at 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow. These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations. Consequently, the solutions derived from these linearized models may not closely align with the actual optimal solutions in real-world scenarios. This trade-off between simplicity and accuracy necessitates careful consideration when applying these linearizations in practice.
In response to these challenges, various studies have explored adaptive power flow approximations tailored to specific systems and operating ranges to enhance approximation accuracy (e.g., optimization-based approaches in [11] and sample-based approaches in [12, 13, 14]); see [15, 16, 17] for recent survey papers on this concept. For sample-based approaches, samples of operating points computed by repeatedly solving the power flow equations at various points within an operating range (e.g., a specified range of power injections) are leveraged to compute the approximations. By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest. These adaptive power flow linearizations spend computational effort in computing the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems. By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
@@
In summary, the main contributions of this paper are:
@@
    Numerical benchmarking of CBLAs for a variety of test cases.
The remainder of this paper is organized as follows: Section II covers background material on the power flow equations as well as sample-based conservative linear approximations. Section III introduces the conservative bias linear approximation approach. Section IV provides numerical results of our approach. Section V concludes the paper along with directions for our future work.
@@ II-A The Power Flow Equations
Consider a power system where a reference bus has the voltage angle set to 00. Let $V(\theta)$ denote the voltage magnitude (phasor). Let $P(Q)$ denote the active (reactive) power injection. We use the subscript $(\cdot)_{i}$ to represent a quantity at bus $i$ and the subscript $(\cdot)_{ik}$ to represent a quantity from or connecting bus $i$ to $k$. Let $j=\sqrt{-1}$. The AC power flow equations at bus $i$ are:
@@
where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G(B)$ is a real (imaginary) part of the admittance matrix associated with the system, and $\mathcal{B}_{i}$ represents the set of all neighboring buses to bus $i$ including bus $i$ itself.
@@ II-B Conservative linear approximations
The nonlinearity of the power flow equations in (1) contribute to the complexity encountered in solving optimization problems. To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig. 1)[12]. Moreover, CLAs facilitate parallel computation by enabling concurrent computation of the CLA for each quantity of interest. The construction of a CLA entails sampling power injections across an operational range of interest, such as a range of loads and power generated by Distributed Energy Resources (DERs), followed by computing power flow solutions for each sample and solving a constrained-regression problem.
For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$. This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }d\in\mathcal{N_{D}}\}$ where $(\,\cdot\,)_{L_{d}}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, and the superscripts max (min) indicate upper (lower) limits.
@@
where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4). This CLA is constructed to ensure the fulfillment of the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
@@ III Conservative Bias Linear Approximations
The CLA approach presented in Section II-B is consistently conservative within the set of drawn samples. However, in certain scenarios, the conservativeness property may lead to significant errors due to specific samples. In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range. The CBLA approach is designed to be optimal, aiming to minimize a specific error metric while retaining a tendency to be conservative in order to enhance accuracy. This implies that the CBLA primarily minimizes errors between the approximating function and the samples, permitting samples to violate conservativeness at a specified cost.
@@ III-A Formulation
The optimization problem in (5) seeks to minimize the aggregated value of the error function $f(\,\cdot\,)$ defined in (7) over all samples by computing the coefficients $a_{0}$ and $\bm{a_{1}}$ in (6). This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$). The error function is computed based on the sign of the error for each sample. In cases of overestimation, the value of $h(\epsilon_{m})$ is designed to be high, imposing a substantial cost for violation. Conversely, the value of $g(\epsilon_{m})$ is intended to be relatively low, reflecting scenarios where samples do not violate the conservativeness. Vice-versa, the value of $h(\epsilon_{m})$ is designed to be relatively low and $g(\epsilon_{m})$ to be high for an underestimation.
@@ III-B Error function
The error function defined in equation (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
@@
In this setup, by assigning a higher penalty when an approximation violates conservativeness, we incline towards more conservative approximations while sacrificing some accuracy. Additionally, when the derivative of the function $h(\epsilon)$ increases as shown in Fig. 3, this function likely permits only small positive values of $\epsilon$, as larger positive values result in exponentially higher penalties. Conversely, reducing the cost tends to yield more accurate approximations at the expense of conservativeness.
@@
With nonlinear error functions, the regression problem is a mixed-integer nonlinear program; nevertheless, it can be conveniently implemented using a user-defined function in Julia and subsequently solved with a Julia package like Optim, which provides a framework for solving constrained optimization problems [18].
IV Numerical results

### v2

# Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA

@@ Abstract
The power flow equations are central to many problems in power system planning, analysis, and control. However, their inherent non-linearity and non-convexity present substantial challenges during problem-solving processes, especially for optimization problems. Accordingly, linear approximations are commonly employed to streamline computations, although this can often entail compromises in accuracy and feasibility. This paper proposes an approach termed Conservative Bias Linear Approximations (CBLA) for addressing these limitations. By minimizing approximation errors across a specified operating range while incorporating conservativeness (over- or under-estimating quantities of interest), CBLA strikes a balance between accuracy and tractability by maintaining linear constraints. By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy. We illustrate the effectiveness of our proposed approach through several test cases, including its application to a unit commitment problem, where CBLA consistently achieves lower operating costs and improved feasibility compared to traditional linearization methods.
@@
Conservative bias linear approximation; power flow approximation; unit commitment.
@@ I Introduction
The power flow equations play a central role in the operation and analysis of electric power systems. These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7]. However, the nonlinearity of the power flow equations induces non-convexities in these problems that pose significant computational challenges.
To address these challenges, researchers have developed various linear approximations such as DC power flow [8], LinDistFlow [9], first-order Taylor expansions of the power flow equations, and other approximations [10]. These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems. However, these linearizations often depend on broad assumptions such as maintaining voltages near 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow. These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations. Consequently, the solutions derived from these linearized models may not closely align with the actual optimal solutions in real-world scenarios. This trade-off between simplicity and accuracy necessitates careful consideration when applying these linearizations in practice.
In response to these challenges, various studies have explored adaptive power flow approximations tailored to specific systems and operating ranges to enhance approximation accuracy (e.g., optimization-based approaches in [11] and sample-based approaches in [12, 13, 14]); see [15, 16, 17] for recent survey papers on this concept. For sample-based approaches, samples of operating points computed by repeatedly solving the power flow equations at various points within an operating range (e.g., a specified range of power injections) are leveraged to compute the approximations. By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest. These adaptive power flow linearizations spend computational effort to calculate the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems. By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
@@
This work extends our previous study presented in [18], where we introduced the CBLA approach for power flow linearization. In this journal version, we build upon that foundation by incorporating a unit commitment formulation [19] using CBLA, comparing it against other power flow linearization methods, and detailing a step-by-step process for solving the unit commitment problem.
In summary, the paper’s main contributions are:
@@
    Numerical tests of CBLAs for a variety of test cases.
  4. (iv)
    An application and numerical results on proposed unit commitment problems comparing the CBLA approach with other power flow linearizations.
The remainder of this paper is organized as follows: Section II presents background material on power flow equations and sample-based conservative linear approximations. Section III introduces the proposed conservative bias linear approximation (CBLA) approach. Section IV describes the iterative method for solving the unit commitment problem using CBLA. Section V provides numerical results demonstrating the performance of the proposed method. Finally, Section VI concludes the paper and outlines directions for future research.
@@ II-A The Power Flow Equations
Consider a power system where a reference bus has the voltage angle set to 00. Let $V(\theta)$ denote the voltage magnitude (angle). Let $P(Q)$ denote the active (reactive) power injection. We use the subscript $(\cdot)_{i}$ to represent a quantity at bus $i$ and the subscript $(\cdot)_{ik}$ to represent a quantity from or connecting bus $i$ to $k$. Let $j=\sqrt{-1}$. The AC power flow equations at bus $i$ are:
@@
where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G$ and $B$ are the real and imaginary parts of the system’s admittance matrix, respectively, and $\mathcal{B}_{i}$ is the set of all neighboring buses to bus $i$, including bus $i$.
@@ II-B Conservative linear approximations
The nonlinearity of the power flow equations in (1) contributes to the complexity encountered in solving optimization problems. To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig. 1)[12]. Moreover, CLAs facilitate parallel computation by enabling concurrent computation of the CLA for each quantity of interest. The construction of a CLA entails sampling power injections across an operational range of interest, such as a range of loads and power generated by Distributed Energy Resources (DERs), followed by computing power flow solutions for each sample and solving a constrained-regression problem.
For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$. This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }L_{d}\in\mathcal{N_{D}}\}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, $\mathcal{N_{D}}$ is the set of all buses with loads, and the superscripts max (min) indicate upper (lower) limits.
@@
where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4). This CLA is designed to ensure the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
@@ III Conservative Bias Linear Approximations
The CLA approach presented in Section II-B is consistently conservative within the set of drawn samples. However, in certain scenarios, the conservativeness property may lead to significant errors due to specific samples. In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range. The CBLA approach is designed to be optimal, minimizing a specific error metric while maintaining a tendency to be conservative to enhance accuracy. This means CBLA focuses on minimizing errors between the approximating function and the samples, allowing some samples to violate conservativeness at a specified cost.
@@ III-A Formulation
The optimization problem in (5) seeks to minimize the aggregated value of the error function $f(\,\cdot\,)$ defined in (7) over all samples by computing the coefficients $a_{0}$ and $\bm{a_{1}}$ in (6). This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$). The error function is computed based on the error’s sign for each sample. For overestimation, $h(\epsilon_{m})$ is set high to impose a substantial cost for violation, while $g(\epsilon_{m})$ is kept low for non-violations. Conversely, for underestimation, $h(\epsilon_{m})$ is relatively low and $g(\epsilon_{m})$ is high. For an alternative formulation of CBLA that incorporates a penalty term to influence the optimization behavior, see the appendix.
@@ III-B Error function
The error function in (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
@@
In this setup, assigning a higher penalty for violations of conservativeness leads to more conservative approximations at the expense of some accuracy. Additionally, when the derivative of the function $h(\epsilon)$ increases, as shown in Fig. 3, the function tends to allow only small positive values of $\epsilon$, because larger positive values result in exponentially higher penalties. Conversely, reducing the penalty cost tends to produce more accurate approximations but with less conservativeness.
@@
Nonlinear error functions transform the regression problem into a mixed-integer nonlinear program. This can be efficiently implemented using a user-defined function in Julia and solved with packages such as Optim, which offers a robust framework for constrained optimization problems [20].
IV Unit Commitment
The unit commitment (UC) problem is an optimization problem in power system operations that aims to determine the optimal schedule of generating units over a specified time horizon, typically 24 hours. The primary goal is to minimize the total cost of operation while ensuring that the demand for electricity is met and the technical and operational constraints of the generation units and transmission network are satisfied.
The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the unit commitment (UC) problem must address to ensure grid reliability and operational efficiency. As a mixed-integer optimization problem, the UC formulation involves binary variables to represent the on/off status of generators, along with continuous variables to model power output.
IV-A Computational procedure and parallelization potential
Figure 4: A flowchart illustrating the computational process for solving the unit commitment (UC) problem using CBLA. Steps enclosed in red boxes correspond to CBLA, while steps enclosed in blue boxes relate to the UC problem. Steps marked with * indicate parallelizable processes.
Solving the unit commitment problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations. Given the complexity of unit commitment and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations. The overall process consists of three key steps:
  1. 1.
    CBLA computation: Solve the CBLA formulation in (5) to approximate voltage and current over a range of operating conditions. This step requires generating samples within expected operating ranges and solving the CBLA optimization problem.
  2. 2.
    Unit commitment optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
  3. 3.
    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [21].
This computational framework allows parallel execution, particularly in Steps 1 and 3. In Step 1, solving CBLA involves two key computations that can be parallelized. First, the power flow solutions at each operating point can be computed independently. Second, the regression problem for the CBLA can be solved independently for each quantity, using the same set of samples across all regression problems. Similarly, in Step 3, since ramping constraints are not enforced and only minimum up- and down-time constraints are considered, the AC OPF computations at different time steps are independent once generator statuses are fixed, enabling parallel execution across time periods. Fig. 4 illustrates the full set of computational steps required to solve the unit commitment problem using the CBLA approach.
IV-B Optimization Problem Formulation
Let $u_{g,t}$ be a generator on/off status and $s_{g,t}\;(r_{g,t})$ be a generator startup (shutdown) status for generator $g$ at time $t$. The unit commitment problem with the DC power flow [22] is formulated as follows:
$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
s.t. $\sum_{g\in\mathcal{G}}P_{g,t}=D_{t},\quad\forall t\in\mathcal{T},$ (10b)
$[\forall g\in\mathcal{G}]$
$P_{g,t}\leq P_{g}^{\text{max}}u_{g,t},\;P_{g,t}\geq P_{g}^{\text{min}}u_{g,t},\;\forall t\in\mathcal{T},$ (10c)
$s_{g,t}-r_{g,t}=u_{g,t}-u_{g,t-1},\;\forall t\in\mathcal{T},$ (10d)
$\sum_{\tau=t}^{t+T_{g}^{\text{up}}-1}u_{g,\tau}\geq T_{g}^{\text{up}}\big{(}u_{g,t}-u_{g,t-1}\big{)},$
$\forall t\in\{1,\ldots,T-T_{g}^{\text{up}}\},$ (10e)
$\sum_{\tau=t}^{t+T_{g}^{\text{down}}-1}(1-u_{g,\tau})\geq T_{g}^{\text{down}}\big{(}u_{g,t-1}-u_{g,t}\big{)},$
$\forall t\in\{1,\ldots,T-T_{g}^{\text{down}}\},$ (10f)
$\theta_{i,t}-\theta_{k,t}=\frac{P_{ik,t}}{B_{ik}},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (10g)
$-P_{ik}^{\text{max}}\leq P_{ik,t}\leq P_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (10h)
where $u_{g,0}=0$ for all $g$, $c_{g}(P_{g,t})$ is a cost associated with generator $g$ at time $t$ (for example, $c_{g}(P_{g,t})=c_{2g}P_{g,t}^{2}+c_{1g}P_{g,t}+c_{0g}$ for a polynomial cost function), $\mathcal{G}$ is a set of all generators, $\mathcal{L}$ is a set of all lines, and $\mathcal{T}$ is a set of time in hours.
The optimization problem in (10) is referred to as the UC problem, which minimizes the total operating cost over a given time horizon. The objective function in (10a) comprises generation costs, startup costs, and shutdown costs.
The constraint (10b) ensures power balance at every time step, requiring total generation to match load demand, where (10c) enforces that the power output of each generator stays within its operational limits when the generator is online. The constraint (10d) models the relationship between the on/off status of generators and their startup/shutdown states. In this setup, we consider ramping constraints as part of the startup and shutdown constraints, ensuring that generation changes smoothly during transitions. The constraints (10e)–(10f) enforce minimum up-time and down-time requirements for generators, ensuring they remain online or offline for a specified duration after being switched on or off. The constraint (10g) represents the linearized DC power flow equations, relating power flows to voltage angle differences and line susceptance. Finally, (10h) ensures that power flow on each line does not exceed its thermal limits in either direction.
To extend this formulation and provide a baseline for comparison, the appendix presents a set of linear constraints based on the first-order Taylor approximation of voltage magnitudes and current flows.
IV-C Unit commitment using conservative bias linear approximations
The unit commitment problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation. While computationally efficient, the DC power flow model does not explicitly incorporate voltage and current constraints. In this section, we introduce penalty terms and replace the DC power flow constraints in (10g)–(10h) with a CBLA, which provides a linearized yet conservative representation of these constraints.
Let $\mathcal{B}$ be a set of all buses. The CBLA enforces voltage magnitude limits at each bus using the following inequalities:
$\underline{a}_{0,i,t}+\underline{\bm{a}}_{1,i,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\geq V_{i}^{\text{min}}-v_{i,t}^{\text{min}},\;\forall i\in\mathcal{B},\forall t\in\mathcal{T},$ (11a)
$\overline{a}_{0,i,t}+\overline{\bm{a}}_{1,i,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\leq V_{i}^{\text{max}}+v_{i,t}^{\text{max}},\;\forall i\in\mathcal{B},\forall t\in\mathcal{T},$ (11b)
where $v_{i,t}^{\text{min}}$ and $v_{i,t}^{\text{max}}$ are nonnegative penalty variables that allow violation of the voltage approximations at bus $i$ at time $t$. Here, $\bm{P}_{t}$ ($\bm{Q}_{t}$) denotes the vector of active (reactive) power injections at time $t$, while $V_{i}^{\text{min}}$ and $V_{i}^{\text{max}}$ are the lower and upper voltage magnitude limits at bus $i$. The coefficients $\underline{a}_{0,i}$ and $\underline{\bm{a}}_{1,i}$ correspond to the underestimating voltage approximation at bus $i$, and $\overline{a}_{0,i,t}$ and $\overline{\bm{a}}_{1,i,t}$ correspond to the overestimating voltage approximation.
Additionally, we enforce current flow limits in transmission lines through the following constraint:
$\overline{b}_{0,ik,t}+\overline{\bm{b}}_{1,ik,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\leq I_{ik}^{\text{max}}+z_{ik,t},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (12)
where $z_{ik,t}$ is a nonnegative penalty variable allowing violation of the current flow approximation on line $(i,k)$ at time $t$, $I_{ik,t}^{\text{max}}$ denotes the maximum allowable current on line $(i,k)$, and $\overline{b}_{0,ik,t}$, $\overline{\bm{b}}_{1,ik,t}$ are coefficients for the overestimating current flow.
Let $F(\bm{P}_{g,t},\bm{s}_{g,t},\bm{r}_{g,t})$ denote the objective function in (10a). With the inclusion of penalty terms, the modified objective becomes:
$F(\bm{P},\bm{s},\bm{r})+\sum_{t\in\mathcal{T}}\left(\lambda\sum_{(i,k)\in\mathcal{L}}z_{ik,t}+\gamma\sum_{i\in\mathcal{B}}(v_{i,t}^{\text{max}}+v_{i,t}^{\text{min}})\right),$ (13)
where $\lambda$ and $\gamma$ are the penalty weights associated with violations of current and voltage magnitude constraints, respectively.
The CBLA method constructs linear constraints from coefficients derived at sampled operating points within a specified range. Before solving the UC problem, we can identify buses and lines that do not exhibit voltage or current violations in the sampled data; constraints for these elements remain inactive and can be excluded, reducing computational complexity without compromising feasibility.
By incorporating these constraints into the UC problem, CBLA ensures generation schedules remain feasible with respect to voltage and line flow limits while maintaining computational tractability. Unlike the standard DC power flow model, CBLA systematically accounts for these limits through conservative yet efficient linear approximations. Its sample-based nature also enables constraint screening, identifying likely binding constraints and allowing flexibility in selecting the error function depending on whether voltage or line flow limits are active.
To maintain the validity of CBLA-based constraints during the UC process, we adopt an iterative approach (Fig. 4). In each iteration, we solve the UC problem with the current CBLA, then check if the generation statuses and dispatch levels match those from a first-order Taylor-based method or previous CBLA iterations. We also verify whether power injections stay within the sampled region used to build the CBLA. If violations occur, we identify the affected time periods and buses or lines, draw new samples around the updated solution, recompute the CBLA, and resolve the UC problem. This process prevents accuracy degradation from extrapolation and cycling between inconsistent UC solutions.
This CBLA-based formulation further allows two practical enhancements. First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservatism. Second, it supports the inclusion of penalty terms that allow for controlled constraint violations, enhancing feasibility in cases where a strictly conservative formulation may be overly restrictive.
IV-D Combinatorial Growth of Valid Generator Statuses
In the unit commitment formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences. Nevertheless, the feasible status space still grows exponentially with the number of time periods, making brute-force enumeration impractical.
To quantify this space, we implement an enumeration algorithm that counts binary vectors of length $\lvert\mathcal{T}\rvert$ satisfying a minimum run length $d=T^{\text{up}}=T^{\text{down}}$ for both states, allowing exceptions for an initial short off-run and final short runs. Due to these exceptions, a clean closed-form expression is difficult to derive. Instead, the number of valid vectors is characterized recursively: sequences are built by appending valid runs (of at least length $d$) to shorter valid sequences, naturally leading to a dynamic programming approach.
Let $f(\lvert\mathcal{T}\rvert,s,\ell)$ denote the number of valid binary sequences of length $\lvert\mathcal{T}\rvert$, ending in state $s\in\{0,1\}$ with a current run of length $\ell$. The recurrence relation is:
$f(\lvert\mathcal{T}\rvert,s,\ell)=\begin{cases}\sum\limits_{k=d}^{\lfloor\nicefrac{{\lvert\mathcal{T}\rvert}}{{2}}\rfloor}f(\lvert\mathcal{T}\rvert-k,1-s,k),\;\text{if }\ell=d,\lvert\mathcal{T}\rvert\geq d\\[4.30554pt] f(\lvert\mathcal{T}\rvert-1,s,\ell-1),\;\text{if }\ell>d\\[4.30554pt] 1,\;\text{if }\lvert\mathcal{T}\rvert=\ell\leq d\text{, one of the exceptions holds}\\[4.30554pt] 0,\;\text{otherwise}\end{cases}$
where the exceptions include:
  • Initial 0’s from position 1 with length less than $d$,
  • Final 0’s or 1’s from position $\lvert\mathcal{T}\rvert-d+2$ onward.
The total number of valid sequences is
$F(\lvert\mathcal{T}\rvert)=\sum_{s\in\{0,1\}}\sum_{\ell=1}^{\lvert\mathcal{T}\rvert}f(\lvert\mathcal{T}\rvert,s,\ell).$
For example, with $\lvert\mathcal{T}\rvert=5$ and $d=2$, applying the recurrence yields 13 valid sequences for a single generator. Therefore, for 6 generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems. This growth illustrates why brute-force enumeration of all valid generator commitment combinations quickly becomes intractable. To manage this complexity, we avoid exhaustive search by solving the unit commitment problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses. The algorithm is explained in more detail in Section IV-A.
IV-E AC Optimal Power Flow with Load Shedding
The AC Optimal Power Flow (AC OPF) problem determines the optimal generation dispatch by solving the full nonlinear AC power flow equations while satisfying network constraints. Using generator commitment statuses from the unit commitment problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour. The multi-period formulation minimizes the total cost across all hours while ensuring power balance and network feasibility at each time. If constraints prevent meeting full demand, load shedding is used as a last resort, with $L_{i,t}$ denoting the load shed at bus $i$ and time $t$. The AC OPF with load shedding is formulated as follows:
$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
s.t. $[\forall i\in\mathcal{B},\forall t\in\mathcal{T},\forall g\in\mathcal{G},\forall(i,k)\in\mathcal{L}]$
$P_{i,t}^{\text{gen}}-P_{i,t}^{\text{load}}+L_{i,t}=$
$\quad\sum_{k\in\mathcal{B}}V_{i,t}V_{k,t}\big{(}G_{ik}\cos\theta_{ik,t}+B_{ik}\sin\theta_{ik,t}\big{)},$ (14b)
$Q_{i,t}^{\text{gen}}-Q_{i,t}^{\text{load}}+\beta_{i}L_{i,t}=$
$\quad\sum_{k\in\mathcal{B}}V_{i,t}V_{k,t}\big{(}G_{ik}\sin\theta_{ik,t}-B_{ik}\cos\theta_{ik,t}\big{)},$ (14c)
$P_{g,t}^{\text{min}}u_{g,t}\leq P_{g,t}\leq P_{g,t}^{\text{max}}u_{g,t},$ (14d)
$Q_{g,t}^{\text{min}}u_{g,t}\leq Q_{g,t}\leq Q_{g,t}^{\text{max}}u_{g,t},$ (14e)
$V_{i,t}^{\text{min}}\leq V_{i,t}\leq V_{i,t}^{\text{max}},$ (14f)
$S_{ik,t}=\sqrt{P_{ik,t}^{2}+Q_{ik,t}^{2}}\leq S_{ik}^{\text{max}},$ (14g)
$0\leq L_{i,t}\leq P_{i,t}^{\text{load}},\quad 0\leq\beta_{i}L_{i,t}\leq Q_{i,t}^{\text{load}},$ (14h)
where $\beta_{i}=\tan\alpha_{i}$, given that $\cos\alpha_{i}$ represents the power factor associated with load $i$. The objective function in (14a) minimizes the total cost of generation over all time periods while penalizing load shedding using a function $C^{\text{shed}}(L_{i,t})$ to discourage demand curtailment unless necessary. The power balance constraints in (14b)–(14c) enforce the nonlinear AC power flow equations for active and reactive power at each bus and time period. The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the unit commitment status $u_{g,t}$. The voltage limits in (14f) ensure that bus voltages remain within safe operating ranges. The thermal constraints in (14g) restrict apparent power flows on transmission lines to their rated capacities. Finally, the load shedding constraints in (14h) ensure that the curtailed demand remains non-negative and does not exceed the original load at any bus and time period. This formulation integrates the unit commitment results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
V Numerical results